# A Regression-Based Approach to the CO2 Airborne Fraction: Enhancing Statistical Precision and Tackling Zero Emissions

### v3

@@ Discussion
Our empirical findings show a slightly higher AF than the consensus estimate of $44\%$ (Canadell et al., 2021, p. 676). The regression-based estimate of the AF, using the 1959-2022 sample of the Global Carbon Project data and controlling for El Niño and volcanic activity, is $47.0\%$ (±$\pm 1.1\%$; $1\sigma$), with a $95\%$ confidence interval of $[44.9\%,49.0\%]$.
@@
When El Niño and volcanic activity are excluded from the analysis, the estimate is $44.8\%$ (±$\pm 1.4\%$; $1\sigma$), which is more in line with the commonly reported results. When we apply the same analysis to two alternative data sets, we obtain slightly higher estimates of the AF than those from the Global Carbon Project (Supplementary Information S2.3). The more recent 1992–2022 subsample yields an AF estimate of approximately $46\%$ (±$\pm 1.0\%$; $1\sigma$) for the Global Carbon Project data and slightly higher estimates for the two alternative data sets (Table 1 and Supplementary Information S2.4). To account for possible measurement error in $G_{t}$ and $E_{t}$, we report Deming regressions (Deming, 1943) in Section S2.5 of the Supplementary Information. The results are in line with the analysis reported so far, and we conclude that measurement error is not driving our results.
@@
To summarize the theoretical findings in our study, we conclude that the ratio-based estimator of the AF suffers from three main shortcomings. First, due to its definition as the ratio of changes in atmospheric concentrations to emissions, means and variances do not exist if zero emissions are possible. While this is of no concern on the historical sample, it is important when analyzing the AF on net-zero emissions scenarios. Studies of the past AF (such as Canadell et al., 2007; Raupach et al., 2008; Le Quéré et al., 2009; Knorr, 2009; Ballantyne et al., 2012; Raupach et al., 2014; Ballantyne et al., 2015; Keenan et al., 2016; Bennedsen et al., 2019; van Marle et al., 2022, among others) are most likely not influenced to any substantial degree by this issue, but studies of future low-emission scenarios are affected (e.g. Pressburger et al., 2023). Second, we require stronger Gaussian assumptions on the distribution of the error process compared to the case of the regression-based estimator when a central limit theorem is invoked to compute confidence intervals and $p$-values based on a Gaussian distribution. Alternative methods such as the bootstrap can also be used for this purpose (Efron and Tibshirani, 1993). Again, studies on historical data are most likely not strongly affected by our findings, as Figure 1d suggests Gaussian residuals. Third, the ratio-based estimator converges to the data-generating AF at a slower rate than the regression-based estimator, even if zero-emissions are ruled out and errors are assumed to be normal. Both estimators converge faster than the common $\sqrt{T}$ rate due to the non-stationarity of the two time series variables emissions ($E_{t}$) and changes in atmospheric concentrations ($G_{t}$) and their cointegration. The ratio-based estimator converges at rate $T$ and the regression-based estimator at rate $T^{3/2}$.
@@
The regression-based estimator that we advocate in this study has standard statistical properties, such as existence of first and second moments, it is defined for zero emissions, and it converges to the data-generating AF at a fast rate. A central limit theorem applies without assuming Gaussianity of the regression error, and confidence levels and $p$-values can be computed in the usual way. Based on theoretical arguments, on a simulation study (Supplementary Information, Section S3), and on a historical sample of yearly data, we have shown that the regression-based estimator exhibits lower estimation uncertainty compared to the ratio-based estimator. Finally, we have argued that the regression-based estimator can readily be generalized to a time-varying AF specification with its estimations by means of the Kalman filter and smoother.
@@
The main advantage of the regression-based estimator for a historical sample analysis is increased precision. In cases of net-zero emission scenarios and of future data analysis with emissions approaching zero, it will remain valid, unlike the ratio-based estimator. To illustrate this advantage, we have simulated trajectories for emissions and changes in atmospheric concentrations over the period 2023–2100 consistent with SSP scenarios using the MAGICC reduced-complexity climate model.
@@ Acknowledgements
Code for replication of all results in the main paper and Supplementary Information can be found at https://github.com/mbennedsen/Regression-Approach-to-CO2-Airborne-Fraction.
@@ A Regression-Based Approach to the CO2 Airborne Fraction:
S3 Simulation study: Performance of the estimators in conditions similar to the data sample 19591959195919592022202220222022
@@
S4 Details on the Kalman filter approach to estimating the time-varying airborne fraction
@@
$\alpha_{t+1}$ $=\alpha_{t}+\eta_{t},\quad\eta_{t}\stackrel{{\scriptstyle iid}}{{\sim}}N(0,\sigma_{\eta}^{2}),$
where $E_{t}$ is treated as a covariate (exogenous variable). Given data on $G_{t}$ we may use the Kalman filter to evaluate the likelihood of the model. The parameters $\sigma_{u}^{2}$ and $\sigma_{\eta}^{2}$ may be estimated by the method of maximum likelihood, and the Kalman smoother provides an estimate of the time-varying parameter $\alpha_{t}$, namely $\hat{\alpha}_{2,t}=\widehat{\mathbb{E}}(\alpha_{t}|\text{data})$, where “data” denotes $(G_{t},E_{t})$ for $t=2022,2023,\ldots,2100$. Similarly, the Kalman smoother provides confidence bands around $\hat{\alpha}_{2,t}$. We refer to Durbin and Koopman (2012) for a textbook treatment of state space models and the Kalman filter and smoother.
S4.1 Application to SSP scenarios 2022-2100
Figures S4S8 present the results from applying the regression based estimator $\hat{\alpha}_{2,t}=\widehat{\mathbb{E}}(\alpha_{t}|\text{data})$ of the time-varying airborne fraction $\alpha_{t}$, obtained from the Kalman smoother, to scenarios SSP1-1.9, SSP2-4.5, SSP3-7.0, SSP4-3.4, and SSP5-8.5, respectively. The SSP scenarios were run using the simple climate model MAGICC (Meinshausen et al., 2011). [^7] The SSP scenarios can be run in MAGICC in a web browser via the link https://live.magicc.org/scenarios/bced417f-0f7f-4bb7-8359-792a0a8b0368/overview. Last accessed June 17, 2024.

### v4

@@ Discussion
[…]
@@
When El Niño and volcanic activity are excluded from the analysis, the estimate is $44.8\%$ (±$\pm 1.4\%$; $1\sigma$), which is more in line with the commonly reported results. When we apply the same analysis to two alternative data sets, we obtain slightly higher estimates of the AF than those from the Global Carbon Project (Supplementary Information S2.3). The more recent 1992–2022 subsample yields an AF estimate of approximately $46\%$ (±$\pm 1.0\%$; $1\sigma$) for the Global Carbon Project data and slightly higher estimates for the two alternative data sets (Table 1 and Supplementary Information S2.4). To account for possible measurement error in $G_{t}$ and $E_{t}$, we report Deming regressions (Deming, 1943) which are in line with the results reported so far (Supplementary Information S2.5). We may conclude that measurement error is not driving our results.
@@
To summarize the theoretical findings in our study, we conclude that the ratio-based estimator of the AF suffers from three main shortcomings. First, due to its definition as the ratio of changes in atmospheric concentrations to emissions, means and variances do not exist if zero emissions are possible. While this is of no concern on the historical sample, it is important when analyzing the AF on net-zero emissions scenarios. Studies of the past AF (such as Canadell et al., 2007; Raupach et al., 2008; Le Quéré et al., 2009; Knorr, 2009; Ballantyne et al., 2012; Raupach et al., 2014; Ballantyne et al., 2015; Keenan et al., 2016; Bennedsen et al., 2019; van Marle et al., 2022, among others) are most likely not influenced to any substantial degree by this issue, but studies of future low-emission scenarios are affected (e.g. Pressburger et al., 2023). Second, we require stronger Gaussian assumptions on the distribution of the error process compared to the case of the regression-based estimator when a central limit theorem is invoked to compute confidence intervals and $p$-values based on a Gaussian distribution. Alternative methods such as the bootstrap can also be used for this purpose (Efron and Tibshirani, 1993). Again, studies on historical data are most likely not strongly affected by our findings, as Figure 1d suggests Gaussian residuals. Third, the ratio-based estimator converges to the data-generating AF at a slower rate than the regression-based estimator, even if zero-emissions are ruled out and errors are assumed to be normal. Both estimators converge faster than the common $\sqrt{T}$ rate due to the non-stationarity of the two yearly time series variables emissions ($E_{t}$) and changes in atmospheric concentrations ($G_{t}$) and their cointegration. The ratio-based estimator converges at rate $T$ and the regression-based estimator at rate $T^{3/2}$.
@@
Our preferred regression-based estimator in this study has standard statistical properties, such as existence of first and second moments, it is defined for zero emissions, and it converges to the data-generating AF at a fast rate. A central limit theorem applies without assuming Gaussianity of the regression error, and confidence levels and $p$-values can be computed in the usual way. Based on theoretical arguments, on a simulation study (Supplementary Information S4), and on a historical sample of yearly data, we have shown that the regression-based estimator exhibits lower estimation uncertainty compared to the ratio-based estimator. Finally, we have argued that the regression-based estimator can readily be generalized to a time-varying AF specification with its estimations by means of the Kalman filter and smoother.
@@
The main advantage of the regression-based estimator for a historical sample analysis is increased precision. In cases of net-zero emission scenarios and of future data analysis with emissions approaching zero, it will remain valid, unlike the ratio-based estimator. To illustrate this advantage, we have simulated trajectories for emissions and changes in atmospheric concentrations over the period 2023–2100 consistent with SSP scenarios using the MAGICC reduced-complexity climate model. We regard this development as a first step and consider the use of climate projections from the Coupled Model Intercomparison Project (CMIP) as next steps in our research agenda.
@@ Acknowledgements
Author contributions
All authors contributed equally to the paper.
Data availability
Data are publicly available and can be found at
@@
MATLAB code for replication of all results in the main paper and Supplementary Information can be found at https://github.com/mbennedsen/Regression-Approach-to-CO2-Airborne-Fraction.
@@ A Regression-based Approach to the CO2 Airborne Fraction:
S3 Cumulative airborne fraction
Another alternative to the ratio-based and regression-based estimators of the AF is the cumulative airborne fraction (CAF) estimator, defined as the ratio of the cumulative changes in atmospheric CO2 concentrations and the cumulative CO2 emissions, given by
$\text{CAF}_{t_{0},t}=\sum_{j=t_{0}}^{t}G_{j}\,\Large{/}\,\sum_{j=t_{0}}^{t}E_{j},$ (S3.1)
where $t_{0}$ denotes the first year of the sample and $t$ is the current year or the last year of the sample. Using the Global Carbon Project data 19591959195919592022202220222022, we can calculate the CAF, yielding
$\text{CAF}_{1959,2022}^{\text{GCP}}=\sum_{j=1959}^{2022}G_{j}\,\Large{/}\,\sum_{j=1959}^{2022}E_{j}=\frac{218.03}{491.03}=44.40\%.$
The two alternative data sets, introduced in Section S2.3, result in
$\text{CAF}_{1959,2022}^{\text{H\&C}}=\sum_{j=1959}^{2022}G_{j}\,\Large{/}\,\sum_{j=1959}^{2022}E_{j}=\frac{218.03}{457.65}=47.64\%,$
and
$\text{CAF}_{1959,2022}^{\text{vMa}}=\sum_{j=1959}^{2022}G_{j}\,\Large{/}\,\sum_{j=1959}^{2022}E_{j}=\frac{218.03}{442.07}=49.32\%.$
Due to its cumulative nature, the CAF is not as useful for studying changes and trends in the AF, when compared to the ratio- and regression-based measures considered in the main paper, which utilize the yearly variables $G_{t}$ and $E_{t}$. To illustrate this point, Figure S3 shows the CAF over the period 2023–2100 using the data from the SSP1-2.6 scenario, i.e. the same data as used in Figure 2 of the main paper. Panel a) of the figure shows the $CAF_{2023,t}$ $t=2023,2024,\ldots,2100$, as defined in Equation (S3.1). As can be seen, the CAF is very smooth and generally adjusts slowly to changes in the behavior of the carbon system.
A possible remedy to the “over-smoothness” is to calculate the CAF over a (smaller) fixed window of length $w\geq 1$. Consider therefore the alternative CAF, defined as
$\text{CAF}_{t_{0},t}(w)=\text{CAF}_{\max\{t_{0}-w+1,t_{0}\},t}=\sum_{j=\max\{t_{0}-w+1,t_{0}\}}^{t}G_{j}\,\Large{/}\,\sum_{j=\max\{t_{0}-w+1,t_{0}\}}^{t}E_{j},$ (S3.2)
for a window length $w\geq 1$. This is the “moving-window” version of the CAF, where each data point is calculated using only the preceding $w$ years (when $t-t_{0}+1<w$, then $t-t_{0}+1$ years are used). For instance, $w=1$ corresponds to the conventional ratio-based AF as considered in the main paper, while $w=78$ is the CAF suggested in Equation (S3.1) above (because there are $78$ yearly data points in the 2023–2100 SSP data). For other values of $w$, but with $781<w<$ 78, the $\text{CAF}(w)$ in (S3.2) is calculated using the cumulative data using shorter windows. This resulting CAF may be able to track the changes and trends in the AF better.
If $w$ is chosen too large, the CAF in (S3.2) will inherit the downsides of the “full-sample” CAF in (S3.1), i.e. it will be slow to adapt. Conversely, in case $w$ is too small, this version of the CAF inherits the downsides of the ratio-based estimator considered in the main paper since it is still defined as a ratio. To illustrate these points, Panels b)–f) of Figure S3 show $\text{CAF}_{2023,t}(w)$, $t=2023,2024,\ldots,2100$, for $w=50,20,10,5,1$, respectively. In particular, when choosing a smaller window size $w$, the CAF is indeed tracking the changes in the behavior of the carbon system more closely, while it also has the effect of making the CAF vulnerable to periods where $\sum E_{t}\approx 0$.
S4 Simulation study: Performance of the estimators in conditions similar to the data sample 19591959195919592022202220222022
@@
S5 Details on the Kalman filter approach to estimating the time-varying airborne fraction
@@
$\alpha_{t+1}$ =$=\begin{cases}\alpha_{t}+\eta_{t}&\text{ if }t+1\neq\tau,\\ 1-\alpha_{t}+\eta_{t}&\text{ if }t+1=\tau,\end{cases},\quad\eta_{t}\stackrel{{\scriptstyle iid}}{{\sim}}N(0,\sigma_{\eta}^{2}),$
where $E_{t}$ is treated as a covariate (exogenous variable) and $\tau$ denotes the year where emissions first turn negative, i.e. $\tau=\inf\{t:E_{t}<0\}$. (We use the convention that if $\{t:E_{t}<0\}$ is empty, then $\tau=\infty$.) This is a random walk model for $\alpha_{t}$, i.e. $\alpha_{t+1}=\alpha_{t}+\eta_{t}$, for all years $t$ except at the year $t+1=\tau$, where the transition equation becomes $\alpha_{t+1}=1-\alpha_{t}+\eta_{t}$. [^7] The reason for specifying the model in a way that reflects $\alpha_{t}$ around one at $t=\tau$ is as follows: When $G_{t}<0$ and $E_{t}>0$ we expect that $\alpha_{t}<0$, which means that the carbon sinks absorb more than is emitted. When emissions turn negative, i.e. $E_{t}<0$, then $\alpha_{t}$ switches to being positive. Because we still expect that the sinks absorb more than what is emitted, we would in fact expect that $\alpha_{t}>1$. This features is obtained by the specification $\alpha_{\tau}=1-\alpha_{\tau-1}+\eta_{\tau}$.
Given the data for $G_{t}$, we may use the Kalman filter to evaluate the likelihood of the model. The parameters $\sigma_{u}^{2}$ and $\sigma_{\eta}^{2}$ may be estimated by the method of maximum likelihood, and the Kalman smoother provides an estimate of the time-varying parameter $\alpha_{t}$, namely $\hat{\alpha}_{2,t}=\widehat{\mathbb{E}}(\alpha_{t}|\text{data})$, where “data” denotes $(G_{t},E_{t})$ for $t=2023,2024,\ldots,2100$. Similarly, the Kalman smoother provides confidence bands around $\hat{\alpha}_{2,t}$. We refer to Durbin and Koopman (2012) for a textbook treatment of state space models and the Kalman filter and smoother.
S5.1 Application to SSP scenarios 2023-2100
Figures S5S9 present the results from applying the regression based estimator $\hat{\alpha}_{2,t}=\widehat{\mathbb{E}}(\alpha_{t}|\text{data})$ of the time-varying airborne fraction $\alpha_{t}$, obtained from the Kalman smoother, to scenarios SSP1-1.9, SSP2-4.5, SSP3-7.0, SSP4-3.4, and SSP5-8.5, respectively. The SSP scenarios were run using the simple climate model MAGICC (Meinshausen et al., 2011). [^8] The SSP scenarios can be run in MAGICC in a web browser via the link https://live.magicc.org/scenarios/bced417f-0f7f-4bb7-8359-792a0a8b0368/overview. Last accessed June 17, 2024.
@@
Figure S3: Cumulative airborne fraction (S3.2) calculated from SSP1-2.6 data over the period 20232023202320232100210021002100. a) $w=78$, corresponding to the CAF in (S3.1); b) $w=50$; c) $w=20$; d) $w=10$; e) $w=5$; f) $w=1$, corresponding to the AF ratio-based estimator, denoted by $\widehat{\alpha}_{1,t}$ in the main paper. See Section S3 for details.